# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

### v1

@@ 6 Conclusions
In this work, we showed that the GM model can provide a unified interpretation of the LHC hints near 95 GeV in the $\gamma\gamma$ and $\tau\tau$ channels with a single light $CP$-even custodial singlet. Implementing the RG-improved positive-definiteness conditions for vacuum stability enlarges the viable parameter space relative to traditional bounded-from-below requirements, while remaining consistent with perturbative unitarity, electroweak precision data, $B$-physics observables, direct searches, and Higgs signal-strength measurements.
The surviving parameter region displays distinctive features that make the scenario predictive: singlet mixing is restricted to narrow intervals; the triplet VEV $v_{\Delta}$ is at the few 10 GeV scale; and the custodial multiplets exhibit a mass hierarchy with $m_{H_{3}}<m_{H_{5}}$ in the ${\cal O}(10^{2})~{\rm GeV}$ order. The $\gamma\gamma$ production rate is enhanced by charged-scalar loops, most notably from the doubly charged state $H_{5}^{\pm\pm}$, which interfere constructively with the SM contributions. In contrast, the $\tau\tau$ signal strength is determined by the universal rescaling of fermionic couplings and the total decay width. Consequently, the scan readily reproduces the reported $2\sigma$ interval for $\mu_{\gamma\gamma}$, while $\mu_{\tau\tau}$ typically remains below the experimental central value but still within approximately $2\sigma$ compatibility.
These features make the framework highly testable in future experiments. At the HL-LHC, updated searches in the $\gamma\gamma$ and $\tau\tau$ final states, along with direct probes of $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$, will be able to explore much of the remaining parameter space. Future lepton colliders, such as CEPC and FCC-ee, will further offer decisive tests of this scenario.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 12447167 and 12275067), by the Joint Fund of Henan Province Science and Technology R&\&D Program (Grant Nos.225200810030, 225200810092 and 225200810014), by the Startup Research Fund of Henan Academy of Sciences (Grant No. 231820011), by the Basic Research Fund of Henan Academy of Sciences (Grant No. 240620006), by the Graduate Innovation Fund of Henan Academy of Sciences (Grant No. 243320031), by the Science and Technology Innovation Leading Talent Support Program of Henan Province, by the National Key R&\&D Program of China (Grant No. 2023YFA1606000) and by the Science and Technology Innovation Leading Talent Support Program of Henan Province (Grant No. 254000510039). PZ is supported by the ARC Discovery Project DP220100007 and by the Centre for the Subatomic Structure of Matter (CSSM).

### v2

@@ 6 Conclusions
In this work, we demonstrated that the GM model offers a unified explanation for the LHC hints near 95 GeV in the $\gamma\gamma$ and $\tau\tau$ channels using a single light $CP$-even custodial singlet. Implementing the RG-improved positive-definiteness conditions for vacuum stability enlarges the viable parameter space relative to traditional bounded-from-below requirements, while remaining consistent with perturbative unitarity, electroweak precision data, $B$-physics observables, direct searches, and Higgs signal-strength measurements.
Our fitting results clearly demonstrate the predictive characteristics of the parameter space: the singlet mixing parameter $\alpha$ is confined to a narrow interval around $\frac{\pi}{3}$; the triplet VEV $v_{\Delta}$ is at the few 10 GeV scale; and the custodial multiplets exhibit a mass hierarchy with $m_{H_{3}}<m_{H_{5}}$ at the ${\cal O}(10^{2})~{\rm GeV}$ scale. The $\gamma\gamma$ production rate is enhanced by charged-scalar loops, most notably from the doubly charged state $H_{5}^{\pm\pm}$, which interfere constructively with the SM contributions. The GM model can naturally accommodate the diphoton excess, but due to the tension from the Higgs data, the di-tau signal strength $\mu_{\tau\tau}$ can only be pushed to around 0.5, which is about $1\sigma$ away from the experimental central value.
These features make our framework highly testable in future experiments. At the HL-LHC, updated searches in the $\gamma\gamma$ and $\tau\tau$ final states, along with direct probes of $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$, will be able to explore much of the remaining parameter space. Future lepton colliders, such as CEPC and FCC-ee, will further offer decisive tests of this scenario.
@@ Acknowledgements.
We thank Cheng-Wei Chiang and Martin J. White for helpful discussions. We acknowledge the referee’s insightful comments. This work was supported by the National Natural Science Foundation of China (Grant Nos. 12447167 and 12275067), by the Joint Fund of Henan Province Science and Technology R&\&D Program (Grant Nos.225200810030, 225200810092 and 225200810014), by the Startup Research Fund of Henan Academy of Sciences (Grant No. 231820011), by the Basic Research Fund of Henan Academy of Sciences (Grant No. 240620006), by the Graduate Innovation Fund of Henan Academy of Sciences (Grant No. 243320031), by the National Key R&\&D Program of China (Grant No. 2023YFA1606000) and by the Science and Technology Innovation Leading Talent Support Program of Henan Province (Grant No. 254000510039). PZ is supported by the ARC Discovery Project DP220100007 and by the Centre for the Subatomic Structure of Matter (CSSM).
Data Availability Statement
All figures presented in this work were generated using the Jarvis-PLOT framework. The complete plotting configuration files used in this study, together with additional result figures not shown in the main text, are publicly available on the webpage:
https://pengxuan-zhu-phys.github.io/Jarvis-Docs/jarvisplot/pub/gm95excess/